OpenStudy (anonymous):

Find the derivative of f(x) = 5 /x at x = -1. 5 1 -1 -5

OpenStudy (jdoe0001):

keep in mind that \(\bf \cfrac{1}{x} \implies x^{-1}\)

OpenStudy (anonymous):

how do i simplify x^-1..? @jdoe0001

OpenStudy (anonymous):

@Ahmad1 ?

OpenStudy (anonymous):

\[f(x)=\frac{ 5 }{ x } = 5(x)^{-1} \] \[f \prime (x) = -5 (x)^{-2} = \frac{ -5 }{ x ^{2} }\] at x=1we get \[\frac{ -5 }{ 1 } = -5\]